find two supplemaentary angles such that the measure of the first angle is four times the measure of the second
Answer by checkley79(3341)

X+4X=180
5X=180
X=180/5
X=36 DEGREES FOR THE SMALLER ANGLE.
180-36=144 DEGREES FOR THE LARGER ANGLE.
PROOF:
5*36=180
180=180